OpenStudy (anonymous):

find the limit as x approaches infinity, -(x+1) (e^(1/(x+1))-1)

OpenStudy (anonymous):

help please

ganeshie8 (ganeshie8):

are you allowed to use L'hops rule ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my problem is how to differentiate those numbers

ganeshie8 (ganeshie8):

\[\large \lim \limits_{x\to \infty }-(x+1) (e^{1/(x+1)}-1) = \lim \limits_{x\to \infty }- \dfrac{ (e^{1/(x+1)}-1)}{1/(x+1)} \]

ganeshie8 (ganeshie8):

substitute \(\large t = 1/(x+1) \) as \(\large x \to \infty\), \(\large t \to 0\) right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

the limit changes to \[\large \lim \limits_{t\to 0 }- \dfrac{ e^{t}-1}{t} \]

ganeshie8 (ganeshie8):

since its of form 0/0 you can apply L'hops

ganeshie8 (ganeshie8):

\[\large \lim \limits_{t\to 0 }- \dfrac{ e^{t}-1}{t} ~\stackrel{LH}{=}~ \lim \limits_{t\to 0 }- \dfrac{ e^{t}}{1} \]

OpenStudy (anonymous):

are we letting t=1/(x+1)?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so after getting l\[\lim_{t \rightarrow 0}\frac{- e ^{t} }{ 1}\] do we substitute t=1/(x+1) back?

ganeshie8 (ganeshie8):

nope, x's are gone the final answer will be a numeral value just pluging t = 0 and evaluate

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

\[\large \lim_{t \rightarrow 0}\frac{- e ^{t} }{ 1} = -e^0 = -1\]

OpenStudy (anonymous):

will the answer be -1?

ganeshie8 (ganeshie8):

thats it!

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

you're welcome!

OpenStudy (anonymous):

so it means anytime you have a question like this, you can represent the exponent by a variable and differentiate.